# Non-invertible duality defects in one, two, and three dimensions via gauging spatially modulated symmetry

### v1

@@ 3.2 Three dimensions
In this subsection, we turn to construction of the duality defects via gauging subsystem symmetries in the 3D Ising plaquette model.
@@
Now we construct non-invertible duality defect operator by gauging subsystem symmetries for spins $\tilde{Z}_{\mathbf{r}}$ [^5] The procedure of gauging such symmetries in 3D was discussed in [5]. To this end, we accommodate the extended Hilbert space on each plaquette with Pauli operators represented by $\tilde{\tau}^{Z}$ and $\tilde{\tau}^{X}$. The Gauss law reads (see also left term in Fig. 2(b))
@@
Here, $\partial\mathbf{p}_{ab}\in\mathbf{r}$ denotes plaquettes on $ab$-plane sharing a corner with a node at $\mathbf{r}$. Plaquette Ising terms $\tilde{P}_{\mathbf{p}_{ab}}$ are minimally coupled to the gauge field as
@@
to ensure that the theory becomes dynamically trivial (See also right configurations in Fig. 2(b).). Here, $\prod_{\mathbf{p}_{ab}\in\partial\mathbf{c}}$ denotes product of the gauge fields over plaquettes on $ab$-plane that surround the cube $\mathbf{c}$. After gauging, we finally arrive at the following gauged Hamiltonian:
$\widehat{H}_{3D:plaquette}=$displaystyle-- $J\sum_{\mathbf{p}_{ab}}P_{Z,\mathbf{p}_{ab}}-h\sum_{\mathbf{r}}X_{\mathbf{r}}-{J}\sum_{\mathbf{p}_{ab}}{\tau^{Z}}_{\mathbf{p}_{ab}}-{h}\sum_{\mathbf{r}}G_{X,\mathbf{r}}$ (46)
-- $g_{x}\sum_{\mathbf{c}}B_{\mathbf{c},x}-g_{y}\sum_{\mathbf{c}}B_{\mathbf{c},y}-g_{z}\sum_{\mathbf{c}}B_{\mathbf{c},z}.$
@@
The duality defect operator is constructed from the swap operator after gauging. Similar to the argument in the previous subsection, a local spin operator $\tilde{Z}_{\mathbf{r}}$ in the swap operator (40) does not commute with the subsystem symmetries, $\tilde{Q}_{ab}$, hence, it should become non-local operators after the gauging. We proceed by replacing $\tilde{Z}_{\mathbf{r}}$ with the Wilson operator $\mathcal{M}_{\mathbf{r}}$ comprised of product of the gauge fields, forming membranes, that is,
@@
Note that $S_{\mathbf{r}}$ with the first (second) relation in (48) is invertible (non-invertible) by noticing that $S_{\mathbf{r}}^{2}=1~{}(S_{\mathbf{r}}^{2}\neq 1)$. The operator $S_{(0,0,0)}$ does not commute with $S_{\mathbf{r}}$ with $\mathbf{r}$ being on $xy$-, $yz$-, $zx$-planes that intersect the origin. Also, $S_{(\hat{x},0,0)}$ does not commute with $S_{\mathbf{r}}$ on $yz$-plane that intersects $x$-axis with the coordinate $(\hat{x},0,0)$, and the similar relation holds when permuting $x$, $y$, and $z$. The operator $S$ is non-invertible due to the fact that $S_{\mathbf{r}}^{2}\neq 1$ in the second case of (48).
@@
As seen from the last two relations, the points at the origin and three spatial axes, the operator $S$ plays the role of the duality defects implementing the first mapping in (43) whereas acting $S$ on the spin around the boundary involves subsystem global charges. To fix the issue, one multiplies subsystem charges with $S$, namely, $S$ is modified as $SC_{Q_{sub}}$, where
@@
$C_{\eta}$ $=$ $\frac{1}{2^{L_{x}+L_{y}+L_{z}-1}}\left[\frac{1}{2}\prod_{\hat{z}=1}^{L_{z}}\left(1+\eta^{xz}_{x}(\hat{z})\right)\left(1+\eta^{zy}_{x}(\hat{z})\right)\right]\times\left[\frac{1}{2}\prod_{\hat{x}=1}^{L_{x}}\left(1+\eta^{xz}_{z}(\hat{x})\right)\left(1+\eta^{xy}_{y}(\hat{x})\right)\right]$ (57)
×$\times\left[\frac{1}{2}\prod_{\hat{y}=1}^{L_{y}}\left(1+\eta^{xy}_{x}(\hat{y})\right)\left(1+\eta^{yz}_{x}(\hat{y})\right)\right],$
@@
$\eta_{a}^{ab}(\hat{b})\vcentcolon=\prod_{\hat{a}=1}^{L_{a}}\tau^{Z}_{\mathbf{p}_{ab}},\quad\eta_{a}^{ab}(\hat{a})\vcentcolon=\prod_{\hat{b}=1}^{L_{b}}\tau^{Z}_{\mathbf{p}_{ab}},$ (58)
which corresponds to noncontractible of the gauge fields, associated with lineon excitations of the $X$-cube model.  [^7] The prefactor $\frac{1}{2}$ inside the bracket in (58) is introduced so that it takes care of the redundancy due to the relation $\prod_{\hat{b}=1}^{L_{b}}\eta^{ab}_{a}(\hat{b})=\prod_{\hat{a}=1}^{L_{a}}\eta^{ab}_{b}(\hat{a})$. The operator in (57) is the subsystem analog of the condensation defects in continuum limit [48] [^8] See also [49, 50] for earlier expositions on the related topic.
To summarize the argument, we have constructed duality defect $D$ (54), satisfying
@@
This duality defect $D$ is non-invertible. Indeed, we obtain the following fusion rule:
$D\times D=C,\quad C\times C=2^{2(L_{x}+L_{y}+L_{z})-4}C,\quad DC=CD=2^{2(L_{x}+L_{y}+L_{z})-4}C.$ (60)
@@
$\begin{split}C=&\left\{\frac{1}{2^{2}}\prod_{\hat{z}=1}^{L_{z}}\left(1+Q_{sub_{xy},\hat{z}}\right)\times\prod_{\hat{x}=1}^{L_{x}}\left(1+Q_{sub_{yz},\hat{x}}\right)\times\prod_{\hat{y}=1}^{L_{y}}\left(1+Q_{sub_{zx},\hat{y}}\right)\right\}\\ &\times\frac{1}{2^{L_{x}+L_{y}+L_{z}-1}}\left[\frac{1}{2}\prod_{\hat{z}=1}^{L_{z}}\left(1+\eta^{xz}_{x}(\hat{z})\right)\left(1+\eta^{zy}_{x}(\hat{z})\right)\right]\times\left[\frac{1}{2}\prod_{\hat{x}=1}^{L_{x}}\left(1+\eta^{xz}_{z}(\hat{x})\right)\left(1+\eta^{xy}_{y}(\hat{x})\right)\right]\\ &\times\left[\frac{1}{2}\prod_{\hat{y}=1}^{L_{y}}\left(1+\eta^{xy}_{x}(\hat{y})\right)\left(1+\eta^{yz}_{x}(\hat{y})\right)\right],\end{split}$ (61)
@@
We conclude this section with a comment that the non-invertible duality defects $D$ for subsystem symmetries discussed in this section and dipole symmetries discussed in Sec. 4 are translational invariant and Hermitian, which follows from similar arguments in Appendix E of Ref. [39].
@@ 4 Dipole symmetry
In this section, we turn to the case of the dipole symmetry in one and two dimensions. As we will demonstrate, the fusion rules of the duality defects are characterized by dipole algebra, which is a new structure that intertwines global and dipole charges.
Before delving into details of the construction, we clarify the usage of the terminology “dipole symmetry”, since it is important in the remaining parts of this work. As we show momentarily, such as (63), we think of the global dipole symmetry, characterized by global charges and dipole charges. Here, we emphasize the word “global” that appears twice in the sentence by italic and teletype fonts, respectively. While the word “global” emphasized by teletype font is used in the usual sense, meaning, the word global is associated with homogeneous spin flip on entire space such as $Q_{0}$ in (63), the usage of the word “global” in the italic font is related to several spin flips including inhomogeneous ones such as $Q_{dipole}$ and $Q_{0}$ in (63). In this section, we consistently use “00-form dipole symmetry” to refer to the global symmetry associated with series of charges including inhomogeneous ones.
@@ 4.1 One dimension
In this subsection, we study the duality defect in the $\mathbb{Z}_{N}$ chain model respecting 00-form dipole symmetry. To do so, we introduce several notations. We define $Z_{j}$ and $X_{j}$ by the clock and shift matrices at site index $j$, respectively. The commutation relation for onsite spins satisfies $Z_{j}X_{j}=\omega X_{j}Z_{j}$, where $\omega=e^{2\pi i/N}$. Also, denoting the system size as $L$, we impose a periodic boundary condition so that $X_{j=0}=X_{j=L}$ and for simplicity, we assume $L=kN\;(k\in\mathbb{Z})$.
@@
$S_{j}X_{i}=\tilde{X}_{j},\quad S_{j}Z_{j}=\tilde{Z}_{j},\quad S_{j}\tilde{X}_{j}=X_{j},\quad S_{j}\tilde{Z}_{j}=Z_{j}.$ (66)
@@
Now we are in a good place to perform gauging the 00-form dipole symmetry $\tilde{Q}_{0}$ and $\tilde{Q}_{dipole}$, viz 00-form dipole symmetry for spins $\tilde{Z}_{i}$ and construct the duality defect. To do so, we introduce an extended Hilbert space located at each node of the chain whose $\mathbb{Z}_{N}$ Pauli operator is denoted as $\tilde{\tau}^{X}_{j}$ and $\tilde{\tau}^{Z}_{j}$, corresponding to gauge fields. The Gauss law is given by
@@
Similar to the discussion presented in Sec. 2, one would expect that the operator $S$ plays the role of the duality defect, sending $X_{j}$ and $Z_{j-1}(Z_{j}^{\dagger})^{2}Z_{j+1}$ to $G_{j}$ and $\tau^{Z}_{j}$, respectively, based on the fact that $S$ transforms a spin operator without tilde into the one with tilde before the gauging and these are further mapped to the operators according to (69) after gauging. To see this is correct, we need to carefully check the form of $S$ after gauging. After gauging the 00-form dipole symmetry, a local spin flip term $\tilde{X}_{j}$ in (65) is replaced by $G_{j}$ due to the Gauss law (67). However, since $\tilde{Z}_{j}$ does not commute with the 00-form dipole symmetry (63), it becomes non-local after gauging, involving multiplication of the gauge fields. To resolve this issue, one replace a local spin $\tilde{Z}_{j}$ with the Wilson operator $W_{j}$ described by nonlocal string of the gauge fields, meaning,
@@
We check whether the operator $S$ after gauging is the duality defect. After some algebra, one finds
$SX_{j}=G_{i}S,\quad SG_{j}=X_{j}S,\quad S\tau^{Z}_{j}=Z_{j-1}(Z_{j}^{\dagger})^{2}Z_{j+1}S,\quad SZ_{j-1}(Z_{j}^{\dagger})^{2}Z_{j+1}=\tau^{Z}_{j}S\quad(2\leq j\leq L-1),$ (76)
hence, away from the point $j=1,L$, the product of the swap operator is the desired duality operator. However, taking care of the periodic boundary condition, we have
@@
This operator is the desired duality defect, which can be verified by
@@
Further, the duality defect is non-invertible; we obtain the following fusion rules:
@@
implying the fusion rule of the duality defects involves dipole and global charges. Note that previous study showed that fusion rule of the non-invertible duality defect obtained by gauging 00-form dipole symmetry in a single chain yields charge conjugation operator in addition to the dipole and global charges. [43, 44]. As opposed to the previous considerations, in our setting comprised of double copies of the spin chains, there is no charge conjugation operator in the fusion rule.

### v2

@@ 3.2 Three dimensions
In this subsection, we turn to construction of the operators via gauging subsystem symmetries in the 3D Ising plaquette model.
@@
Now we construct noninvertible operator operator by gauging subsystem symmetries for spins $\tilde{Z}_{\mathbf{r}}$ [^7] The procedure of gauging such symmetries in 3D was discussed in [5]. To this end, we accommodate the extended Hilbert space on each plaquette with Pauli operators represented by $\tilde{\tau}^{Z}$ and $\tilde{\tau}^{X}$. The Gauss law reads (see also left term in Fig. 2(b))
@@
Here, $\partial\mathbf{p}_{ab}\in\mathbf{r}$ denotes plaquettes on $ab$-plane sharing a corner with a node at $\mathbf{r}$. Similar to the previous subsection, intuitive understanding of the Gauss law term (46) is that one crops the global charges into local ones; defining
$\mathcal{G}_{sub3D,\mathbf{r}}\vcentcolon=\tilde{X}_{\mathbf{r}}\times\prod_{\partial\mathbf{p}_{xy}\in\mathbf{r}}\tilde{\tau}^{X}_{\mathbf{p}_{xy}}\times\prod_{\partial\mathbf{p}_{yz}\in\mathbf{r}}\tilde{\tau}^{X}_{\mathbf{p}_{yz}}\times\prod_{\partial\mathbf{p}_{zx}\in\mathbf{r}}\tilde{\tau}^{X}_{\mathbf{p}_{zx}},$
it follows that
$Q_{sub_{ab},\hat{c}}=\prod_{\hat{a}=1}^{L_{a}}\prod_{\hat{b}=1}^{L_{b}}\mathcal{G}_{sub3D,\mathbf{r}}\quad(1\leq\hat{c}\leq L_{c}),$ (47)
where $a,b,c$ are cyclic permutations of $x,y,z$. The Gauss law is imposed by setting $\mathcal{G}_{sub3D,\mathbf{r}}=1\quad\forall\mathbf{r}$.
Plaquette Ising terms $\tilde{P}_{\mathbf{p}_{ab}}$ are minimally coupled to the gauge field as
@@
to ensure that the theory becomes dynamically trivial (See also right configurations in Fig. 2(b).) which amounts to imposing flatness condition of the gauge fields on the theory. Here, $\prod_{\mathbf{p}_{ab}\in\partial\mathbf{c}}$ denotes product of the gauge fields over plaquettes on $ab$-plane that surround the cube $\mathbf{c}$. After gauging, we finally arrive at the following gauged Hamiltonian:
$\widehat{H}_{3D:plaquette}=$displaystyle-- $J\sum_{\mathbf{p}_{ab}}P_{Z,\mathbf{p}_{ab}}-h\sum_{\mathbf{r}}X_{\mathbf{r}}-{J}\sum_{\mathbf{p}_{ab}}{\tau^{Z}}_{\mathbf{p}_{ab}}$ (52)
-- ${h}\sum_{\mathbf{r}}G_{X,\mathbf{r}}-g_{x}\sum_{\mathbf{c}}B_{\mathbf{c},x}-g_{y}\sum_{\mathbf{c}}B_{\mathbf{c},y}-g_{z}\sum_{\mathbf{c}}B_{\mathbf{c},z}.$
@@
The operator is constructed from the swap operator after gauging. Similar to the argument in the previous subsection, a local spin operator $\tilde{Z}_{\mathbf{r}}$ in the swap operator (45) does not commute with the subsystem symmetries, $\tilde{Q}_{ab}$, hence, it should become non-local operators after the gauging. We proceed by replacing $\tilde{Z}_{\mathbf{r}}$ with the Wilson operator $\mathcal{M}_{\mathbf{r}}$ comprised of product of the gauge fields, forming membranes, that is,
@@
Note that $S_{\mathbf{r}}$ with the first (second) relation in (54) is invertible (noninvertible) by noticing that $S_{\mathbf{r}}^{2}=1~{}(S_{\mathbf{r}}^{2}\neq 1)$. The operator $S_{(0,0,0)}$ does not commute with $S_{\mathbf{r}}$ with $\mathbf{r}$ being on $xy$-, $yz$-, $zx$-planes that intersect the origin. Also, $S_{(\hat{x},0,0)}$ does not commute with $S_{\mathbf{r}}$ on $yz$-plane that intersects $x$-axis with the coordinate $(\hat{x},0,0)$, and the similar relation holds when permuting $x$, $y$, and $z$. The operator $S$ is noninvertible due to the fact that $S_{\mathbf{r}}^{2}\neq 1$ in the second case of (54).
@@
As seen from the last two relations, the points at the origin and three spatial axes, the operator $S$ plays the role of the operators implementing the first mapping in (49) whereas acting $S$ on the spin around the boundary involves subsystem global charges. To fix the issue, one multiplies subsystem charges with $S$, namely, $S$ is modified as $SC_{Q_{sub}}$, where
@@
$C_{\eta}$ $=$ $\frac{1}{2^{L_{x}+L_{y}+L_{z}-1}}\left[\frac{1}{2}\prod_{\hat{z}=1}^{L_{z}}\left(1+\eta^{zx}_{x}(\hat{z})\right)\left(1+\eta^{yz}_{y}(\hat{z})\right)\right]\times\left[\frac{1}{2}\prod_{\hat{x}=1}^{L_{x}}\left(1+\eta^{zx}_{z}(\hat{x})\right)\left(1+\eta^{xy}_{y}(\hat{x})\right)\right]$ (63)
×$\times\left[\frac{1}{2}\prod_{\hat{y}=1}^{L_{y}}\left(1+\eta^{xy}_{x}(\hat{y})\right)\left(1+\eta^{yz}_{z}(\hat{y})\right)\right],$
@@
$\eta_{a}^{ab}(\hat{b})\vcentcolon=\prod_{\hat{a}=1}^{L_{a}}\tau^{Z}_{\mathbf{p}_{ab}},\quad\eta_{b}^{ab}(\hat{a})\vcentcolon=\prod_{\hat{b}=1}^{L_{b}}\tau^{Z}_{\mathbf{p}_{ab}},$ (64)
which corresponds to noncontractible of the gauge fields, associated with lineon excitations of the $X$-cube model.  [^9] The prefactor $\frac{1}{2}$ inside the bracket in (63) is introduced so that it takes care of the redundancy due to the relation $\prod_{\hat{b}=1}^{L_{b}}\eta^{ab}_{a}(\hat{b})=\prod_{\hat{a}=1}^{L_{a}}\eta^{ab}_{b}(\hat{a})$. As an example, $\eta^{xy}_{x}(\hat{y})$ [$\eta^{xy}_{y}(\hat{x})$] denotes a non-contractible loop of the lineon in the $x$[$y$]-direction formed by gauge fields on a $xy$-plane at the coordinate $\hat{y}[\hat{x}]$. Note that these loops do not have $\hat{z}$ dependence since they are topological in the $z$-direction. The operator in (63) is the subsystem analog of the condensation defects in continuum limit [55] [^10] See also [56, 57] for earlier expositions on the related topic.
To summarize the argument, we have constructed operator $D$ (60), satisfying
@@
This operator $D$ is noninvertible. Indeed, we obtain the following fusion rule:
$D\times D=C,\quad C\times C=2^{2(L_{x}+L_{y}+L_{z})-4}C,\quad DC=CD=2^{2(L_{x}+L_{y}+L_{z})-4}D.$ (66)
@@
$\begin{split}C=&\left\{\frac{1}{2^{2}}\prod_{\hat{z}=1}^{L_{z}}\left(1+Q_{sub_{xy},\hat{z}}\right)\times\prod_{\hat{x}=1}^{L_{x}}\left(1+Q_{sub_{yz},\hat{x}}\right)\times\prod_{\hat{y}=1}^{L_{y}}\left(1+Q_{sub_{zx},\hat{y}}\right)\right\}\\ &\times\frac{1}{2^{L_{x}+L_{y}+L_{z}-1}}\left[\frac{1}{2}\prod_{\hat{z}=1}^{L_{z}}\left(1+\eta^{zx}_{x}(\hat{z})\right)\left(1+\eta^{yz}_{y}(\hat{z})\right)\right]\times\left[\frac{1}{2}\prod_{\hat{x}=1}^{L_{x}}\left(1+\eta^{zx}_{z}(\hat{x})\right)\left(1+\eta^{xy}_{y}(\hat{x})\right)\right]\\ &\times\left[\frac{1}{2}\prod_{\hat{y}=1}^{L_{y}}\left(1+\eta^{xy}_{x}(\hat{y})\right)\left(1+\eta^{yz}_{z}(\hat{y})\right)\right],\end{split}$ (67)
@@
We conclude this section with a comment that the noninvertible operators $D$ for subsystem symmetries discussed in this section and dipole symmetries discussed in Sec. 4 are translational invariant and Hermitian, which follows from similar arguments in Appendix E of Ref. [42].
@@ 4 Dipole symmetry
In this section, we turn to the case of the dipole symmetry in one and two dimensions. As we will demonstrate, the fusion rules of the operators are characterized by dipole algebra, which is a new structure that intertwines global and dipole charges. In what follows, we consistently use “00-form dipole symmetry” to refer to the two $\mathbb{Z}_{N}$ generators as global $\mathbb{Z}_{N}$ uniform and dipole symmetries.
@@ 4.1 One dimension
In this subsection, we study the operator in the $\mathbb{Z}_{N}$ chain model respecting 00-form dipole symmetry. To do so, we introduce several notations. We define $Z_{j}$ and $X_{j}$ by the clock and shift matrices at site index $j$, respectively. The commutation relation for onsite spins satisfies $Z_{j}X_{j}=\omega X_{j}Z_{j}$, where $\omega=e^{2\pi i/N}$. Also, denoting the system size as $L$, we impose a periodic boundary condition so that $X_{j=0}=X_{j=L}$ and for simplicity, we assume $L=kN\;(k\in\mathbb{Z})$.
@@
$S_{j}X_{j}=\tilde{X}_{j},\quad S_{j}Z_{j}=\tilde{Z}_{j},\quad S_{j}\tilde{X}_{j}=X_{j},\quad S_{j}\tilde{Z}_{j}=Z_{j}.$ (72)
@@
Now we are in a good place to perform gauging the 00-form dipole symmetry $\tilde{Q}_{0}$ and $\tilde{Q}_{dipole}$, viz 00-form dipole symmetry for spins $\tilde{Z}_{i}$ and construct the operator. To do so, we introduce an extended Hilbert space located at each node of the chain whose $\mathbb{Z}_{N}$ Pauli operator is denoted as $\tilde{\tau}^{X}_{j}$ and $\tilde{\tau}^{Z}_{j}$, corresponding to gauge fields. The Gauss law is given by
@@
Intuitive understanding of the Gauss law term (73) is that one decomposes the global charges (69) into local ones by introducing extended Hilbert spaces associated with the gauge fields. To see how, we define
$\mathcal{G}_{1Ddip,j}\vcentcolon=(\tilde{\tau}^{X}_{j-1})^{\dagger}\tilde{\tau}^{X}_{j}\tilde{X}_{j}\tilde{\tau}^{X}_{j}(\tilde{\tau}^{X}_{j+1})^{\dagger}$
and obtain the following:
$\tilde{Q}_{0}=\prod_{j=1}^{L}\mathcal{G}_{1Ddip,j},\quad\tilde{Q}_{dipole}=\prod_{j=1}^{L}(\mathcal{G}_{1Ddip,j})^{j}.$ (74)
The Gauss law is imposed by setting $\mathcal{G}_{1Ddip,j}=1\quad\forall j$.
@@
Similar to the discussion presented in Sec. 2, one would expect that the operator $S$ plays the role of the operator, sending $X_{j}$ and $Z_{j-1}(Z_{j}^{\dagger})^{2}Z_{j+1}$ to $G_{j}$ and $\tau^{Z}_{j}$, respectively, based on the fact that $S$ transforms a spin operator without tilde into the one with tilde before the gauging and these are further mapped to the operators according to (76) after gauging. To see this is correct, we need to carefully check the form of $S$ after gauging. After gauging the 00-form dipole symmetry, a local spin flip term $\tilde{X}_{j}$ in (71) is replaced by $G_{j}$ due to the Gauss law (73). However, since $\tilde{Z}_{j}$ does not commute with the 00-form dipole symmetry (69), it becomes non-local after gauging, involving multiplication of the gauge fields. To resolve this issue, one replace a local spin $\tilde{Z}_{j}$ with the Wilson operator $W_{j}$ described by nonlocal string of the gauge fields, meaning,
@@
Note that operators $\{S_{j}|1\leq j\leq L-2\}$ commute with one another whereas $S_{L-1}$ does not with $S_{j}(j\neq 0\mod N)$. Also, $S_{L}$ does not commute with $S_{j}(j\neq 1\mod N)$.
@@
We check whether the operator $S$ after gauging is the operator. After some algebra, one finds
$SX_{j}=G_{j}S,\quad SG_{j}=X_{j}S,\quad S\tau^{Z}_{j}=Z_{j-1}(Z_{j}^{\dagger})^{2}Z_{j+1}S,\quad SZ_{j-1}(Z_{j}^{\dagger})^{2}Z_{j+1}=\tau^{Z}_{j}S\quad(2\leq j\leq L-1),$ (83)
hence, away from the point $j=1,L$, the product of the swap operator is the desired operator. However, taking care of the periodic boundary condition, we have
@@
This operator is the desired swap operator, which can be verified by
@@
Further, the operator is noninvertible; we obtain the following fusion rules:
@@
implying the fusion rule of the operators involves dipole and global charges. Note that previous study showed that fusion rule of the noninvertible operator obtained by gauging 00-form dipole symmetry in a single chain yields charge conjugation operator in addition to the dipole and global charges. [47, 48]. As opposed to the previous considerations, in our setting comprised of double copies of the spin chains, there is no charge conjugation operator in the fusion rule.